# dApps: Enabling Real-Time AI-Based Open RAN Control

### v1

@@ Abstract
Open RAN , dApps , Real-Time Control Loops , Radio Resource Management (RRM) , Spectrum Sharing , Positioning
@@ 1 Introduction
How to bring AI/ML to O-RAN and push inference toward the real-time domain has received increasing interest in the last few years. Apart from dApps, the literature offers different proposals that include the concepts of real-time RIC and $\mu$Apps [16], zApps [17], tApps [18], to name a few. In this direction, we recently edited an O-RAN nGRG research report [19], with input and collaborations across academia and industry aimed at exploring use cases and applications that would benefit from dApps. These include real-time scheduler reconfiguration, spectrum sensing, compute resource scaling for energy savings, channel equalization, beam management and many others. Despite the well-defined use cases, an architecture for implementing dApps in cellular networks is still lacking. In particular, a comprehensive description of the procedures, interfaces, modules, and their interactions with existing O-RAN components—such as xApps, rApps, RICs, DUs, CUs, and open interfaces (e.g., O1, O2, and E2)—is still missing.
@@
Additionally, we implement two distinct dApp use cases using the proposed framework: positioning and spectrum sharing. The results reveal that the positioning dApp is able to compute the distance between the User Equipment (UE) and the Next Generation Node Base (gNB) using the UE Uplink (UL) Channel Impulse Response (CIR) collected by the gNB in real-time, and with the advantage of plug-and-play, customizable processing routines. Finally, the spectrum-sharing dApp effectively detects incumbents in the 5th generation (5G) gNB RF context, enabling more efficient spectrum utilization through real-time analysis.
@@
The remainder of this paper is organized as follows. Section 2 describes the role of the dApps within the O-RAN architecture and the use cases that can benefit from their implementation. Section 3 presents the integration of the dApps in the O-RAN architecture, introducing the E3 logic interface and describing the message exchange between the dApps and the other O-RAN components. Section 4 describes the lifecycle of the dApps, from their onboarding through the Service Management and Orchestration (SMO) to their deployment and in the RAN unit. In Section 5, we provide a reference implementation of the architecture proposed based on OAI and a custom Python framework to measure the real-time control loop communication. We also discuss benchmarking results that profile the performance of the control loop, demonstrating that our real-time control loop can execute in less than $450\,\mu s$. In Section 6, we use our framework to develop two dApp-powered use cases: Spectrum Sharing and Sensing and Positioning. We evaluate their performance and impact on the RAN, and show how dApps can detect incumbents in a 5G network and enable spectrum sharing with them. Section 7 presents the current state of the art for the real-time control loops in O-RAN, highlighting the differences between our work and the real-time RICs proposed in the literature. Section 8 concludes the paper and envisions future developments for dApps. Further clarifications on the dApp deployment process are discussed in A.
@@ 2 The Role of dApps in the Hierarchical O-RAN Control Architecture
Fig. 1 reports a simplified view of the O-RAN logical architecture and control loops. This includes the RAN, with the Open CUs-Control Plane (CP) and User Plane (UP), the DU, and the Radio Unit (RU). [^2] For the purposes of this paper, we consistently refer to the Open RAN units while omitting the prefix “O-”. The F1 interface connects CU and DU, while the fronthaul interface bridges the DU and the RU. The E1 interface is in between the CU-CP and UP network functions. The O-RAN architecture leverages the Non-RT RIC and the SMO framework for orchestration and policy definition, through the O1 interface to all RAN nodes and the A1 interface to the Near-RT RICs. The latter performs radio resource management with direct control of the RAN nodes through the E2 interface. These elements are deployed over a set of physical and virtualized resources on a heterogeneous infrastructure (i.e., the O-Cloud). The O2 interface from the SMO interacts with the O-Cloud to configure the compute layer.
@@
However, O-RAN xApp/rApp-based control loops are not suitable for real-time control. For example, transferring I/Q samples or, in general, user data out of the RAN, is generally unfeasible due to security, privacy, and timing/bandwidth constraints, as also mentioned in Sec 1. We discussed this in [15, 19], estimating that transferring I/Q samples through a rate-limited E2 interface would take seconds, which is incompatible with real-time interactions. Bringing programmability and observability to RAN nodes is, thus, paramount to address both limitations. The bottom part of Fig. 1 illustrates the role that dApps can play in future iterations of the O-RAN architecture. dApps are lightweight and plug-and-play services that provide secure and real-time access to the stack without introducing the overhead that an additional real-time RIC would bring to network-oriented RAN nodes.
@@
    Direct Processing of Waveforms and PDUs. Access to I/Q samples can be used for physical layer security, e.g., through anomaly detection to detect spoofing of base stations or legitimate users [10], and with RF fingerprinting for secure authentication. Similarly, information carried by I/Q samples on scheduled or unscheduled symbols can be used to detect spectrum holes or incumbents. This enables spectrum sharing but also remote interference detection, including from other base stations in the operator network. In both cases, information processed at a dApp level can be provided to the stack or, for example, xApps that can coordinate the response across multiple RAN nodes.
@@
    Real-Time Scheduling and Beam Management. This includes real-time scheduling reconfiguration to embed ad hoc policies and Service Level Agreements (SLAs) for specific traffic patterns and profiles (e.g., new slices), scheduling acceleration for efficient massive Multiple Input, Multiple Output (MIMO) and multi-cell coordination [28], and real-time sharing coordination with arbitration across different entities [29]. Similarly, beam management procedures can be enhanced and aided by direct access and manipulation of reference and synchronization signals in the DU [30].
@@
Based on the use cases and related requirements discussed in Sec. 2, in the following paragraphs we discuss how the O-RAN architecture can be extended to support dApps as custom, pluggable RAN microservices. We focus on three key elements: (i) the integration of dApps with RAN nodes, extended via a service-based E3 Application Programming Interface (API); (ii) the proposed E3 API procedures that enable seamless message exchange between the RAN and the dApps; and (iii) the interactions between dApps and xApps, facilitated through our E3 interface and the O-RAN E2 interface to the Near-RT RIC. Further discussion on the dApps lifecycle and the role of the O1 interface is provided in Sec. 4. Finally, Sec. 5 describes a prototype based on this architecture and the open-source OAI stack as a reference.
@@
The DU RAN function can provide the dApp with access to a variety of information elements in the data plane. I/Q samples can be accessed pre- or post-equalization, depending on the dApp requirements, at different levels of granularity and periodicity. Such data can be collected from various channels such as Physical Uplink Control Channel (PUCCH) or Physical Uplink Shared Channel (PUSCH), and can include all subcarriers or specific subsets. The Buffer Status Reports (BSRs), exchanged by Medium Access Control (MAC) and RLC instances, provide real-time information about buffer conditions and can be streamed at different intervals (e.g., per slot or subframe) or accessed on-demand. Channel Quality Information (CQI) values are generated based on O-DU processing and can be streamed immediately or polled on demand to support dynamic resource allocation. CSI can be similarly streamed in real-time or accessed on-demand for adaptive modulation and coding decisions. Finally, uplink Sounding Reference Signal (SRS), transmitted by the UE and processed by the O-DU, can be made available either continuously or on request, providing insights for uplink resource configuration.
@@ 3.2 Service-Based dApps E3 API Endpoint
To streamline dApp operations, the E3 agent consolidates access to RAN data, such as telemetry or data-plane units (e.g., I/Q samples). A centralized subscription manager within the agent handles all message dispatching between the dApps and the RAN. This unified access mechanism eliminates redundant protocol stack calls and ensures efficient coordination by allowing dApps to register callbacks for querying control and data functionalities. The design simplifies interactions and maintains low latency, supporting seamless integration with RAN operations without introducing overhead. The E3 APIs enable dApps to extract necessary RAN data and deliver computed control actions or inferences without disrupting RAN operations. The API design ensures concurrent functionality across multiple dApps while maintaining operational integrity. Given the co-location of CUs/DUs and dApps on the same host, the E3 API leverages Inter-Process Communication (IPC) mechanisms—such as shared memory, logical First In First Out (FIFO) queues, function calls, or system-domain sockets—instead of network PDUs or sockets. This approach is demonstrated in our prototype implementation (Sec. 5).
@@ 3.4 E2SM-DAPP: a service model for managing interactions between dApps and xApps
This structured LCM approach is divided into three main phases: (i) Development, in which applications are designed and built; (ii) Onboarding, in which applications are integrated into the target environment; and (iii) Operations, in which applications are managed and maintained in production [36]. Fig. 5 provides a high-level representation of dApp LCM, with particular focus on the onboarding and operations phases. While the development process is common to all types of applications (i.e., xApps, rApps, and dApps), Fig. 5 focuses specifically on unique aspects related to onboarding and operation of dApps. In the following, we provide a detailed overview of how these phases can be extended and adapted to dApp LCM.
@@ 5 A Reference Open-Source dApp Framework
The ZMQConnector class uses ZeroMQ as the link layer, an high-performance asynchronous messaging library, aimed at use in distributed or concurrent applications [38]. ZeroMQ abstracts the complexity of operations such as connection management and data serialization via simple APIs calls, and it is designed for both intra- and inter-machine communication with flexible built-in patterns, as discusses below. As illustrated in Figure 7, our approach leverages ZeroMQ. Specifically, the initial E3 setup employs a request-reply (REQ-REP) pattern—where the dApp issues the request (REQ) and the RAN side provides the reply (REP)—since this procedure occurs only once during the dApp’s lifecycle. In contrast, subsequent outbound and inbound data exchanges adopt a publish-subscribe (PUB-SUB) pattern, with the data-generating entity acting as the publisher (PUB) and the data-processing entity as the subscriber (SUB). This choice reflects the need to support continuous data transfer between the RAN and the dApp, in contrast to the one-time setup operation.
@@
Data received through the E3Connector socket, regardless of the underlying link and transport layers, is processed by a callback mechanism operated by the subscription manager. This mechanism associates newly available data with the registered dApps that require it, thus enabling real-time data handling and avoiding data duplication. For the PDUs used for establishing communications between the RAN node and dApps, we employ Abstract Syntax Notation One (ASN.1) definitions for message formatting in the E3AP and E3SM, adhering to established standard 3rd Generation Partnership Project (3GPP) and O-RAN message exchange protocols. ASN.1 enables efficient, rapid, and compact serialization of data. Each class extending our DApp abstract class should implement its own SM and intelligent logic according to its use case, as shown in Fig. 6.
@@ 5.3 Evaluating real-time capabilities of dApps
Collect DataProcess DataCreate ControlDeliver ControlCumulativeOperations0100200300400Microseconds [$\mu$s]Control[bytes]01632128
@@
Collect DataProcess DataCreate ControlDeliver ControlCumulativeOperations0100200300400Microseconds [$\mu$s]Control[bytes]01632256
@@
Collect DataProcess DataCreate ControlDeliver ControlCumulativeOperations0100200300400Microseconds [$\mu$s]Control[bytes]01632208
@@
Collect DataProcess DataCreate ControlDeliver ControlCumulativeOperations0100200300400Microseconds [$\mu$s]Control[bytes]01632380
@@
Results in Fig. 8 demonstrate that our dApp prototype achieves real-time control loops with an average aggregated latency of 400 $\mu$s, consistently remaining below 450 $\mu$s. As expected, the usage of IPC through ZeroMQ substantially reduces the impact of the socket operations on the total control loop time, making the operation of Data Collection the indication message and Control Delivery the fastest operations in the loop. The major contribution to the elapsing time of the loop is represented by the Elaboration of the Indication Message, i.e., the ASN.1 decode and the intelligent control logic implemented by the dApp to analyze the data. This behavior is also highlighted in the case of Control Messages with size 64 and 95 bytes, shown in Figs. 8(b) and 8(d), respectively, where the presence of padding in the E3AP PDU causes additional memory allocation, thus increasing the elaboration time. The average total elapsed time for each of the 16 cases is consistently below the 1 ms threshold as shown in Fig. 8, which is defined as the limit for real-time use cases involving the manipulation of I/Q samples (see Fig. 1, Sec. 2).

### v2

@@ Abstract
Open RAN , dApps , Real-Time Control Loops , Radio Resource Management (RRM) , Spectrum Sharing , Positioning , Integrated Sensing and Communication (ISAC)
@@ 1 Introduction
[…]
@@
Additionally, we implement two distinct dApp use cases using the proposed framework for Integrated Sensing and Communication (ISAC): positioning and spectrum sharing. The results reveal that the positioning dApp is able to compute the distance between the User Equipment (UE) and the Next Generation Node Base (gNB) using the UE Uplink (UL) Channel Impulse Response (CIR) collected by the gNB in real-time, and with the advantage of plug-and-play, customizable processing routines. Finally, the spectrum-sharing dApp effectively detects incumbents in the 5th generation (5G) gNB RF context, enabling more efficient spectrum utilization through real-time analysis.
[…]
@@
The remainder of this paper is organized as follows. Section 2 presents the current state of the art for the real-time control loops in O-RAN, highlighting the differences between our work and the real-time RICs proposed in the literature. Section 3 describes the role of the dApps within the O-RAN architecture and the use cases that can benefit from their implementation. Section 4 presents the integration of the dApps in the O-RAN architecture, introducing the E3 logic interface and describing the message exchange between the dApps and the other O-RAN components. Section 5 describes the lifecycle of the dApps, from their onboarding through the Service Management and Orchestration (SMO) to their deployment and in the RAN unit. In Section 6, we provide a reference implementation of the architecture proposed based on OAI and a custom Python framework to measure the real-time control loop communication. We also discuss benchmarking results that profile the performance of the control loop, demonstrating that our real-time control loop can execute in less than $450\,\mu s$. In Section 7, we use our framework to develop two dApp-powered use cases: Spectrum Sharing and Sensing and Positioning. We evaluate their performance and impact on the RAN, and show how dApps can detect incumbents in a 5G network and enable spectrum sharing with them. Section 8 concludes the paper and envisions future developments for dApps. Further clarifications on the dApp deployment process are discussed in A.
2 Related Work and Comparison with Real-time RIC
The need for the introduction of real-time control loops in O-RAN has been advocated in recent works to control lower-layer functionalities of the gNB protocol stack [22, 23]. In our previous work [15], we introduced the concept of dApps to process data locally available at the gNB, and the need for sub-$10$ ms control loops. However, our prior work primarily focused on advocating the benefits of such tighter control loops for generic use cases. In this work, instead, we propose a complete reference architecture and design for dApps, including interfaces for interoperability of dApps with the rest of the O-RAN components. Then, we showcase the capabilities and potential of dApps for use cases of interest to Open RAN networks.
[…]
(a) RT-RIC deployment
(b) dApp deployment
Figure 1: Comparison between RT RIC and dApp architectures. From [19].
While the above-mentioned applications require a dedicated RT RIC entity, as well as additional layers of network abstractions, dApps execute at the CU/DU directly. This minimizes the impact on the E2 nodes and takes advantage of interfaces and functionalities already defined in the O-RAN specifications.
Finally, [24] proposes Janus, a monitoring and control framework that can load and execute custom real-time “codelets” on the CU and DU. Even though this work—possibly the most similar to ours—extends O-RAN to real-time control, Janus’s codelets require direct access to the CU/DU protocol stack rather than interfacing with it, as dApps instead do. In some cases, this can be a limitation as vendors might need to host third-party codelets rather than exposing parameters via interfaces to regulate access to CU/DU functionalities. Moreover, this approach also requires the definition and deployment of a dedicated controller on the RIC, instead of relying on standardized components for the management of real-time RAN functions, as we instead do in our dApp architecture.
Table 1 summarizes the key differences between dApps and other real-time control solutions and the advantages of our proposal.
[…]
Connectivity to RAN stack Disaggregated Disaggregated Integrated
[…]
Deployment Location Directly on CU/DU Require dedicated RT RIC platform Directly modifies CU/DU protocol stack
Additional Network Complexity Minimal, leverages existing RIC components High, add new RIC layers High, modifies CU/DU stack
Resource Utilization Optimized, avoids additional overhead Increased due to extra control layers High, requires additional controllers
Integration with O-RAN Use standardized interfaces and components Require additional network abstractions Limited, modifies protocol stack instead of interfacing with it
Scalability High, built on existing RAN elements Limited, require new infrastructure Limited, protocol stack modifications reduce compatibility
Modularity High, they interface with CU/DU without modifying stack Medium, dependent on RT RIC architecture Low, tightly coupled with CU/DU stack
Flexibility Adapt to different CU/DU implementations Constrained by RT RIC architecture Low, as it relies on direct modifications to protocol stack
Use of Standard O-RAN APIs Yes, integrates with O-RAN standard components Partially, introduce custom abstractions No, modifies stack implementation directly
Table 1: Comparison of dApps with RT RIC-based solutions and Janus.
@@ 3 The Role of dApps in the Hierarchical O-RAN Control Architecture
Figure 2 reports a simplified view of the O-RAN logical architecture and control loops. This includes the RAN, with the Open CUs-Control Plane (CP) and User Plane (UP), the DU, and the Radio Unit (RU). [^2] For the purposes of this paper, we consistently refer to the Open RAN units while omitting the prefix “O-”. The F1 interface connects CU and DU, while the fronthaul interface bridges the DU and the RU. The E1 interface is in between the CU-CP and UP network functions. The O-RAN architecture leverages the Non-RT RIC and the SMO framework for orchestration and policy definition, through the O1 interface to all RAN nodes and the A1 interface to the Near-RT RICs. The latter performs radio resource management with direct control of the RAN nodes through the E2 interface. These elements are deployed over a set of physical and virtualized resources on a heterogeneous infrastructure (i.e., the O-Cloud). The O2 interface from the SMO interacts with the O-Cloud to configure the compute layer.
@@
However, O-RAN xApp/rApp-based control loops are not suitable for real-time control. For example, transferring I/Q samples or, in general, user data out of the RAN, is generally unfeasible due to security, privacy, and timing/bandwidth constraints, as also mentioned in Sec 1. We discussed this in [15, 19], estimating that transferring I/Q samples through a rate-limited E2 interface would take seconds, which is incompatible with real-time interactions, and can limit the capabilities of the E2SM for the Lower Layer Control (LLC) [30]. Bringing programmability and observability to RAN nodes is, thus, paramount to address both limitations. The bottom part of Fig. 2 illustrates the role that dApps can play in future iterations of the O-RAN architecture. dApps are lightweight and plug-and-play services that provide secure and real-time access to the stack without introducing the overhead that an additional real-time RIC would bring to network-oriented RAN nodes.
@@
    Direct Processing of Waveforms and PDUs. Access to I/Q samples can be used for physical layer security, e.g., through anomaly detection to detect spoofing of base stations or legitimate users [10], and with RF fingerprinting for secure authentication. Similarly, information carried by I/Q samples on scheduled or unscheduled symbols can be used to detect spectrum holes or incumbents. This enables spectrum sharing but also remote interference detection, including from other base stations in the operator network. In both cases, information processed at the dApp level can be shared with the stack or xApps to coordinate responses across multiple RAN nodes, paving the way for the RAN to function as an ISAC system.
@@
    Real-Time Scheduling and Beam Management. This includes real-time scheduling reconfiguration to embed ad hoc policies and Service Level Agreements (SLAs) for specific traffic patterns and profiles (e.g., new slices), scheduling acceleration for efficient massive Multiple Input, Multiple Output (MIMO) and multi-cell coordination [31], and real-time sharing coordination with arbitration across different entities [32]. Similarly, beam management procedures can be enhanced and aided by direct access and manipulation of reference and synchronization signals in the DU [33]. In this context, dApps can extract real-time measurements, perform data-driven inference at near-real-time periodicity, and send the resulting data to an xApp, which uses it to control beamforming via the existing E2SM LLC [30].
@@
Based on the use cases and related requirements discussed in Sec. 3, in the following paragraphs we discuss how the O-RAN architecture can be extended to support dApps as custom, pluggable RAN microservices. We focus on three key elements: (i) the integration of dApps with RAN nodes, extended via a service-based E3 interface; (ii) the proposed E3 procedures that enable seamless message exchange between the RAN and the dApps; and (iii) the interactions between dApps and xApps, facilitated through our E3 interface and the O-RAN E2 interface to the Near-RT RIC. Moreover, the E3 interface requires two main components, the E3 API and E3 Agent. The E3 Application Programming Interface (API) is the software implementation of the E3 logical interface, providing the necessary functions for interaction. Meanwhile, the E3 Agent is a entity that implements the E3 interface using the E3 API to facilitate and manage communication. Further discussion on the dApps lifecycle and the role of the O1 interface is provided in Sec. 5. Finally, Sec. 6 describes a prototype of E3 Agent based on this architecture and the open-source OAI stack as a reference.
@@
The DU RAN function can provide the dApp with access to a variety of information elements in the data plane. I/Q samples can be accessed pre- or post-equalization, depending on the dApp requirements, at different levels of granularity and periodicity. Such data can be collected from various channels such as Physical Uplink Control Channel (PUCCH) or Physical Uplink Shared Channel (PUSCH), and can include all subcarriers or specific subsets. The Buffer Status Reports (BSRs), exchanged by MAC and RLC instances, provide real-time information about buffer conditions and can be streamed at different intervals (e.g., per slot or subframe) or accessed on-demand. Channel Quality Information (CQI) values are generated based on O-DU processing and can be streamed immediately or polled on demand to support dynamic resource allocation. CSI can be similarly streamed in real-time or accessed on-demand for adaptive modulation and coding decisions. Finally, uplink Sounding Reference Signal (SRS), transmitted by the UE and processed by the O-DU, can be made available either continuously or on request, providing insights for uplink resource configuration.
@@ 4.2 Service-Based dApps E3 API Endpoint
To streamline dApp operations, the E3 agent consolidates access to RAN data, such as telemetry or data-plane units (e.g., I/Q samples). A centralized subscription manager within the agent handles the dispatching of the management messages between the dApps and the RAN, i.e., the initial pairing, the subscription request and response, and the eventual de-registration of the dApp. This unified access mechanism eliminates redundant protocol stack calls and ensures efficient coordination by allowing dApps to register callbacks for querying control and data functionalities. The design simplifies interactions and maintains low latency, supporting seamless integration with RAN operations without introducing overhead. With The E3 APIs enable dApps to extract necessary RAN data and deliver computed control actions or inferences without disrupting RAN operations. The API design ensures concurrent functionality across multiple dApps while maintaining operational integrity. Given the co-location of CUs/DUs and dApps on the same host, the E3 API leverages Inter-Process Communication (IPC) mechanisms—such as shared memory, logical First In First Out (FIFO) queues, function calls, or system-domain sockets—instead of network PDUs or sockets. This approach is demonstrated in our prototype implementation (Sec. 6).
@@ 4.4 E2SM-DAPP: a service model for managing interactions between dApps and xApps
This structured LCM approach is divided into three main phases: (i) Development, in which applications are designed and built; (ii) Onboarding, in which applications are integrated into the target environment; and (iii) Operations, in which applications are managed and maintained in production [39]. Fig. 6 provides a high-level representation of dApp LCM, with particular focus on the onboarding and operations phases. While the development process is common to all types of applications (i.e., xApps, rApps, and dApps), Fig. 6 focuses specifically on unique aspects related to onboarding and operation of dApps. While the SMO is responsible for deploying the dApp, once deployed, the dApp can operate independently or be managed by other entities, such as an xApp, through the E2SM-DAPP. In the following, we provide a detailed overview of how these phases can be extended and adapted to dApp LCM.
@@ 6 A Reference Open-Source dApp Framework
The ZMQConnector class uses ZeroMQ as the link layer, an high-performance asynchronous messaging library, aimed at use in distributed or concurrent applications [41]. ZeroMQ abstracts the complexity of operations such as connection management and data serialization via simple APIs calls, and it is designed for both intra- and inter-machine communication with flexible built-in patterns, as discusses below. As illustrated in Fig. 8, our approach leverages ZeroMQ. Specifically, the initial E3 setup employs a request-reply (REQ-REP) pattern—where the dApp issues the request (REQ) and the RAN side provides the reply (REP)—since this procedure occurs only once during the dApp’s lifecycle. In contrast, subsequent outbound and inbound data exchanges adopt a publish-subscribe (PUB-SUB) pattern, with the data-generating entity acting as the publisher (PUB) and the data-processing entity as the subscriber (SUB). This choice reflects the need to support continuous data transfer between the RAN and the dApp, in contrast to the one-time setup operation.
@@
Data received through the E3Connector socket, regardless of the underlying link and transport layers, is processed by a callback mechanism operated by the subscription manager. This mechanism associates newly available data with the registered dApps that require it, thus enabling real-time data handling and avoiding data duplication. For the PDUs used for establishing communications between the RAN node and dApps, we employ Abstract Syntax Notation One (ASN.1) definitions for message formatting in the E3AP and E3SM, adhering to established standard 3GPP and O-RAN message exchange protocols. ASN.1 enables efficient, rapid, and compact serialization of data. Each class extending our DApp abstract class should implement its own SM and intelligent logic according to its use case, as shown in Fig. 7.
@@ 6.3 Evaluating real-time capabilities of dApps
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]